Suppose that the marginal rate of substitution is 2, the price of X is sh 3,and the price of Y is sh.1 a. If the consumer obtains 1 more unit of X, how many units of Y must be given up in order to keep utility constant? b. If the consumer obtains 1 more unit of Y, how units many of X must be given up in order to keep utility constant? c. What is the rate at which the consumer is willing to substitute X for Y? d. What is the rate at which the consumer is able to substitute X for Y ?


1
Expert's answer
2022-02-01T10:14:43-0500

"MRS=\frac{Mu_x}{Mu_y}=2"

a) "\frac{-\delta Y}{\delta X}=2"


"-\delta Y= 1 \times 2=2"

"\therefore= \Delta Y= -2"


b) "\frac{-\delta Y}{\delta X}=2"


"\delta X= \frac{2}{-1}=-2"

"\therefore= \Delta X= -2"

c) The consumer is willing to substitute 2 units of X for good Y

d) The consumer is willing to substitute 2 units of Y for X